Let g(x) = 8x + sin 8x for 0 ≤ x < 2π.
The function g will have ____critical points in [0,2pi)
List in increasing order the first two critical points in [0, 2π). Enter DNE for any empty answer blank.
x =
x =
This problem is driving me bonkers! I guess I am confused with how the domain works with this . I got to cos(8x) =-1 part and then did a U-sub to find out what 8x is, I know cos(u)=-1 then u must be pi. I got pi/8 but I am confused beyond belief on how to find the increasing order of this or what it means. Like how it gets to be 3pi/8, 5pi/8 etc.
Thank you for any and all help!

Answer by jim_thompson5910(35256)

g(x) = 8x + sin(8x)
g ' (x) = 8 + 8*cos(8x)
0 = 8 + 8*cos(8x)
8*cos(8x) = -8
cos(8x) = -8/8
cos(8x) = -1
8x = arccos(-1)
8x = pi + 2pi*n ... where n is an integer
x = (pi + 2pi*n)/8


If we made a table with n as the input, x as the output, then...

nx (exact)x (approximate)
-1-pi/8-0.39270
0pi/80.39270
13pi/81.17810
25pi/81.96350
37pi/82.74889
49pi/83.53429
511pi/84.31969
613pi/85.10509
715pi/85.89049
817pi/86.67588



When n = -1, the value of x is -0.39270 which is outside the interval [0,2pi). The same applies to when n = 8 (x = 6.67588). So we only focus on when n = 0 all the way through n = 7.

We can see that there are 8 critical points (7-0+1 = 8) since there are 8 x values in the interval [0,2pi).

The first two critical points are x = pi/8 and x = 3pi/8
pi/8 = 0.39270 approx
3pi/8 = 1.17810 approx
This corresponds to n = 0 and n = 1 respectively.

========================================================================================
Answer:

The function g will have     8     critical points in [0,2pi)
List in increasing order the first two critical points in [0, 2pi). Enter DNE for any empty answer blank.
x = pi/8
x = 3pi/8

If the book is seeking approximate x values (instead of exact), then the two answers for x are
x = 0.39270
x = 1.17810